# Supernova Remnants with Mirror Dark Matter and Hyperons

### v1

@@ I Introduction
Given that the exact nature of DM and its in-medium properties – such as self-interaction and coupling with standard model particles – remains unknown, there are two main approaches to studying DM-admixed neutron stars. One approach involves considering non-gravitational interactions between ordinary matter (OM) and DM, such as the Higgs portal mechanism [4, 5, 6, 7, 8, 9], and self-interacting DM models [10, 11]. In this scenario, the system can be effectively treated as a single fluid due to the non-gravitational interaction between ordinary matter and DM. The other approach involves ignoring the non-gravitational interaction between OM and DM, resulting in a two-fluid system where the different sectors interact solely through gravity. The two-fluid approach has been extensively studied (see e.g. Refs. [12, 13, 14, 15, 6, 16, 17, 18, 19, 20, 21, 22]).
@@
It has also been argued that DM can be captured within a star if it loses its kinetic energy through scattering with nuclear matter in the star. When this process takes place over some time it can lead to the accumulation of DM inside the star [37, 38, 39, 40, 41, 42]. The quantity of DM that can be trapped inside the compact object will depend on the properties of the DM, the type of compact object (NS, white dwarf, or a black hole), and its evolutionary history. For example, WIMPs can accumulate in an NS due to elastic scattering with nucleons. Since the density of OM in the NSs dominates, the energy lost by the DM during collisions can be significant. This enhances the capture of DM inside the compact object [37, 43, 44].
@@
This paper investigates the effect of DM on the evolution of supernova remnants from core birth as a neutrino-rich proto-neutron star (PNS) to maturity as neutrino-poor cold-catalyzed NS [55, 56, 57, 58, 59]. We use the quasi-static approximation, which allows us to assume a spherically symmetric PNS in hydrostatic equilibrium. The stellar evolution is then examined through the evolution of the intensive thermodynamic properties, such as entropy density and lepton fractions, over the Kelvin-Helmholtz timescale [60, 56, 61]. The equations of state of DM and OM are calculated separately under the same thermodynamic conditions, assuming that DM and OM have the same entropy. The effect of DM on the macroscopic and microscopic properties of the PNS is then investigated using the two-fluid formalism, in which DM and nuclear matter interact only gravitationally. This formalism allows us to introduce a fixed DM mass fraction into stellar matter and investigate its effect.
The OM is composed of nucleons, hyperons, and leptons while the DM, on the other hand, comprises self-interacting dark fermions of the same kind. We use the relativistic mean-field approximation adjusted by the density-dependent meson-nucleon couplings (DDME2) [62] for OM and DM species at $\beta$-equilibrium. The coupling for the hyperons is taken from  [63], which extends the DDME2 to include hyperons and $\Delta$-resonances couplings based on symmetry group arguments. he main focus of this work is to examine how the presence of DM influences temperature fluctuations, particle distributions, and sound speed within the star throughout its evolution, assuming gravitational interactions. The structure of the resulting star is modeled, and its tidal deformability is calculated to compare the results with observable constraints.
@@ II.1 Nuclear Matter Model
From the matter EoS, i.e. the $P_{t}$ vs. $\varepsilon_{t}$ relation, one can obtain the free energy relation, $\mathcal{F}_{B}=\varepsilon_{t}-Ts$, which allows us to obtain the entropy density $s$ (where $s=s_{B}n_{B}$, with $s_{B}$ the total entropy of all the particles per baryon and $n_{B}$ the total baryon density). The quantities $P_{t}$ and $\varepsilon_{t}$ are the respective total pressure and total energy density of the matter, encompassing all the baryons and leptons present in the system. For stellar matter, it reads
$sT=\varepsilon_{t}+P_{t}-\sum_{b}\mu_{b}n_{b}-\sum_{L}\mu_{L}n_{L},$ (7)
@@
$sT=P_{t}+\varepsilon_{t}-n_{B}\mu_{B}-\mu_{\nu_{e}}(n_{\nu_{e}}+n_{e}).$ (9)
Equations (8) and (9) are the simplified versions of the $s_{B}$ expression using $\beta$-equilibrium and charge neutrality conditions. We calculate the temperature fluctuations along with the EoS by fixing $s_{B}$ together with the lepton fraction ($Y_{L,e}$) to control the neutrino concentration in the neutrino-trapped stage and $Y_{\nu_{e}}=0$ in the neutrino-free stage. A detailed calculation of the EoS can be found in [73] where we adopted the EoS for this work. In addition to that, density-dependent meson-nucleon coupling has been used to study PNSs with exotic baryons in [74, 55, 58, 75] where EoSs have been explicitly derived.
@@ II.2 Dark Matter Model
${\cal L}_{\rm DM}={}$ ${\overline{\psi}_{D}}[(i\gamma_{\mu}\partial^{\mu}-\gamma^{0}g_{v}V_{0})-(m_{D}-g_{\tilde{\sigma}}\tilde{\sigma}_{0})]\psi_{D}+$
$-\frac{1}{2}m_{\rm\tilde{\sigma}}^{2}\tilde{\sigma}_{0}^{2}+\frac{1}{2}m_{v}^{2}V_{0}^{2},$ (10)
@@
This Lagrangian density is similar to the Walecka model for nucleon (see Ref. [76] and references therein), where we have omitted the isospin projection that differentiates between proton and neutron, thus, referring to them as identical dark fermions. To calculate the EoS for the model, we use the mean-field approximation with density-dependent coupling similar to the approach presented in Sec. II.1 for the OM. Therefore, we replaced $g_{\tilde{\sigma}}$ and $g_{v}$ with $g_{\sigma N}$ and $g_{\omega N}$ and the meson masses $m_{\rm\tilde{\sigma}}$ and $m_{v}$ with $m_{\sigma}$ and $m_{\omega}$ respectively, following Tab. 2, and $m_{D}=939$ MeV. This formalism mirrors the visible sector by adopting the same saturation density, coupling constants, and meson masses as nuclear matter. This assumption reduces the arbitrariness in fixing the couplings. The coupling is defined in the same form as Eq. (5), with $\eta\rightarrow n_{D}/n_{0}$, using the fit parameters on Tab. 2. The detailed calculation of the EoS of the above Lagrangian density at $T=0$ can be found in [31].
@@
$\varepsilon_{D}={}$ $\gamma_{D}\int\frac{d^{3}k_{D}}{(2\pi)^{3}}E_{D}\left[f_{D+}+f_{D-}\right]+\frac{g_{\rm\omega N}^{2}}{2m_{\rm\omega}^{2}}n_{D}^{2}$
@@
$n_{D}=\gamma_{D}\int\frac{d^{3}k_{D}}{(2\pi)^{3}}\left[f_{D\,+}-f_{D\,-}\right],$ (16)
@@
$n_{D}^{s}=\gamma_{D}\int\frac{d^{3}k_{D}}{(2\pi)^{3}}\frac{m^{\ast}_{D}}{E_{D}}\left[f_{D\,+}+f_{D\,-}\right].$ (17)
@@
$\mu_{D}^{\ast}=\mu_{D}-g_{\omega N}V_{0}-\Sigma^{r}_{D},$ (19)
where $\mu_{D}$ is the real chemical potential of the DM and $T_{D}$ is temperature. We calculate the EoS at a fixed dark entropy $\tilde{s}$ ($\tilde{s}=s_{D}n_{D}$, where $\tilde{s}$ is DM entropy density, $s_{D}$ is the total entropy per DM particle) using the free energy expression $\mathcal{F}_{D}=\varepsilon_{D}-T_{D}\tilde{s}$ for dark particles. Similarly to the visible sector, the dark sector is treated as a single fermionic fluid governed by the Pauli exclusion principle. Therefore, a balance between dark fermions and dark leptons must be maintained through a dark $\beta$-equilibrium reaction to accurately characterize the physical properties of the dark stellar system. Since we have identical dark fermions, in the neutrino-trapped matter
$\mu_{D}=\mu_{e^{\prime}}-\mu_{\nu_{e}^{\prime}},$ (20)
where $\mu_{e^{\prime}}$ is the chemical potential of the dark electrons and $\mu_{\nu_{e}^{\prime}}$ is the chemical potential of the dark lepton neutrinos and in the neutrino transparent matter
$\mu_{D}=\mu_{e^{\prime}}.$ (21)
The dark muons become relevant when the matter is neutrino-transparent as mentioned above, however, their chemical potential is the same as that of electrons ($\mu_{e^{\prime}}=\mu_{\mu^{\prime}}$). In the mirrored model scenario, DM in isolation can form compact objects (a dark NS, perhaps), with a simpler composition, assuming the presence of only dark nucleons. Since the dark NSs are physically observable objects in the dark world, we impose dark charge neutrality like the visible sector:
$n_{D}=n_{e^{\prime}},$ (22)
@@
$\tilde{s}T_{D}=P_{D}+\varepsilon_{D}-2n_{D}\mu_{D},$ (24)
@@
$\tilde{s}T_{D}=P_{D}+\varepsilon_{D}-2n_{D}\mu_{D}-\mu_{\nu_{e}^{\prime}}(n_{\nu_{e}^{\prime}}+n_{e^{\prime}}),$ (25)
for neutrino-trapped matter. Studies related to mirror dark matter models and their applications to compact astrophysical objects can be found in [77, 78, 79, 80].
@@ III Two Fluid Formalism
$\frac{dP_{t}}{dr}={}$ $-(P_{t}+\varepsilon_{t})\frac{4\pi r^{3}(P_{t}+P_{D})+M(r)}{r(r-2M(r))},$ (26)
$\frac{dP_{D}}{dr}={}$ $-(P_{D}+\varepsilon_{D})\frac{4\pi r^{3}(P_{t}+P_{D})+M(r)}{r(r-2M(r))},$ (27)
@@
$\frac{dM(r)}{dr}=4\pi(\varepsilon_{t}+\varepsilon_{D})r^{2}.$ (28)
@@
$F(r)=\frac{r-4\pi r^{3}\left\{[\varepsilon_{t}(r)+\varepsilon_{D}(r)]-[P_{t}(r)+P_{D}(r)]\right\}}{r-2M(r)},$ (32)
@@
$Q(r)={}$ $\frac{4\pi r\left\{5[\varepsilon_{t}(r)+\varepsilon_{D}(r)]+9[P_{t}(r)+P_{D}(r)]+\frac{\varepsilon_{t}(r)+P_{t}(r)}{\partial P_{t}(r)/\partial\varepsilon_{t}(r)}+\frac{\varepsilon_{D}(r)+P_{D}(r)}{\partial P_{D}(r)/\partial\varepsilon_{D}(r)}-\frac{6}{4\pi r^{2}}\right\}}{r-2M(r)}$
$-4\left\{\frac{M(r)+4\pi r^{3}[P_{t}(r)+P_{D}(r)]}{r^{2}\left(1-2M(r)/r\right)}\right\}^{2},$ (33)
@@ IV Results
Table 3: Maximum stellar mass ($M_{\rm max}$), its radius ($R_{\rm max}$), the radii of the canonical and $\rm M=2.1M_{\odot}$ stars ($\rm R_{1.4}$ and $\rm R_{2.1}$), and the canonical star tidal deformability ($\Lambda_{1.4}$) for different configurations, OM compositions and proportions of DM mass fraction in an NS.
Configurarion Composition DM [%] $M_{\rm max}$ [$\rm M_{\odot}$] $R_{\rm max}$ [km] $R_{1.4}$ [km] $R_{2.1}$ [km] $\Lambda_{1.4}$
$s_{B}=1$; $Y_{L,e}=0.4$ N 0 2.43 12.49 15.41 14.15 865
1 2.40 12.40 15.37 14.03 704
5 2.30 12.12 14.85 13.47 480
NH 0 2.32 12.53 15.41 14.04 863
1 2.29 12.46 15.37 13.15 708
5 2.19 12.12 14.86 13.88 483
$s_{B}=2$; $Y_{L,e}=0.2$ N 0 2.48 13.01 16.62 15.03 1281
1 2.45 12.93 16.57 14.89 903
5 2.35 12.64 15.96 14.50 596
NH 0 2.30 12.72 16.43 14.42 863
1 2.28 12.65 16.36 14.24 708
5 2.18 12.31 15.74 13.08 483
$s_{B}=2$; $Y_{\nu_{e}}=0.0$ N 0 2.49 13.07 16.97 15.18 1299
1 2.46 13.02 16.95 15.05 1036
5 2.34 12.71 16.31 13.82 753
NH 0 2.28 12.70 16.67 14.36 1152
1 2.25 12.65 16.63 14.15 988
5 2.14 12.33 15.96 13.13 737
$T=0$ N 0 2.48 12.06 13.19 13.20 704
1 2.45 11.94 13.11 13.08 500
5 2.33 11.66 12.76 12.61 356
NH 0 2.26 11.94 13.19 12.98 704
1 2.22 11.82 13.12 12.80 490
5 2.11 11.48 12.75 11.83 348
Figure 1: EoS for dark matter configurations with and without trapped neutrinos (top), and the corresponding temperature fluctuations (bottom).
@@
$i$ represents the different particles in the stellar matter and $n$ is their number densities. Fig. 2 shows the particle distributions within a star with mass $2.1M_{\odot}$ at different stages of the stellar evolution. Along the panels, we present snapshots of the stellar evolution from core birth to the stages where the star heats up by neutrino diffusion and then to ‘maturity’ as cold-catalyzed NS several years after the supernova explosion. The first panel (top-left) represents the ambient conditions of core birth when the star is neutrino-rich. The second panel (top-right) represents the period of deleptonization after the supernova explosion, at this stage the star is about 0.5 to 1 second old. Here, the stellar matter heats up due to neutrino diffusion so its temperature is higher than the first stage. The third stage (bottom-left) occurs when neutrinos fully escape from the stellar core. At this stage, the star reaches its maximum temperature and then begins to cool, roughly 10 to 15 seconds after the core birth. The last panel (bottom-right) is when the star has cooled down through neutrino diffusion and thermal radiation to form a cold-catalyzed NS, over several years.
It is important to mention that two fates await the PNSs after the supernova explosion: If the pressure generated by the supernova explosion is insufficient to completely eject the outer mantle of the progenitor massive star, the PNS may continue to accrete matter, potentially leading to the formation of a black hole. In this case, it is believed that the neutrino emission can abruptly stop [82] and the evolution of the PNS terminates without moving to the second stage. On the other hand, if the pressure generated by the explosion is sufficient to lift the surrounding stellar envelope and induce deleptonization, the outer mantle collapses, allowing the PNS to continue its evolution. During deleptonization, transitioning from the second to the third panel, the likelihood of strangeness-rich particles emerging in the stellar matter increases. This can abruptly soften the EoS, potentially forming a black hole if the gravitational pressure is sufficiently strong [59]. Therefore, Fig. 2 is constructed under the assumption that the PNS continues to evolve without collapsing into a black hole.

### v2

@@ I Introduction
Given that the exact nature of DM and its in-medium properties – such as self-interaction and coupling with standard model particles – remains unknown, there are two main approaches to studying DM-admixed neutron stars. One approach involves considering non-gravitational interactions between ordinary matter (OM) and DM, such as the Higgs portal mechanism [4, 5, 6, 7, 8, 9], and self-interacting DM models [10, 11]. In this scenario, the system can be effectively treated as a single fluid due to the non-gravitational interaction between ordinary matter and DM. The other approach involves ignoring the non-gravitational interaction between OM and DM, resulting in a two-fluid system where the different sectors interact solely through gravity. The two-fluid approach has been extensively studied (see e.g. Refs. [12, 13, 14, 15, 6, 16, 17, 18, 19, 20, 21, 22, 23, 24]).
@@
It has also been argued that DM can be captured within a star if it loses its kinetic energy through scattering with nuclear matter in the star. When this process takes place over some time it can lead to the accumulation of DM inside the star [39, 40, 41, 42, 43, 44]. The quantity of DM that can be trapped inside the compact object will depend on the properties of the DM, the type of compact object (NS or white dwarf), and its evolutionary history. For example, WIMPs can accumulate in an NS due to elastic scattering with nucleons. Since the density of OM in the NSs dominates, the energy lost by the DM during collisions can be significant. This enhances the capture of DM inside the compact object [39, 45, 46].
@@
This paper investigates the effect of DM on the evolution of supernova remnants from core birth as a neutrino-rich proto-neutron star (PNS) to maturity as neutrino-poor, cold-catalyzed NS [57, 58, 59, 60, 61]. We use the quasi-static approximation, which allows us to assume a spherically symmetric PNS in hydrostatic equilibrium. The stellar evolution is then examined through the evolution of the intensive thermodynamic properties, such as entropy density and lepton fractions, over the Kelvin-Helmholtz timescale [62, 58, 63]. The equations of state of DM and OM are calculated separately under the same thermodynamic conditions, assuming that DM and OM have the same entropy and lepton fraction.
However, to the best of our knowledge, at the time of writing this paper, there is no established reference on how DM thermalizes with OM in a way that would justify thermal equilibrium between them. Nonetheless, since we are considering a mirrored fermionic DM model capable of forming isolated dark NSs, we adopt the same entropy per baryon as a reasonable approximation. Given that the interaction between the dark and visible sectors is purely gravitational, thermal equilibrium between the two components is unlikely. Nevertheless, by analogy with the visible sector, where NSs are expected to establish a well-defined entropy distribution, we extend this assumption to the dark sector as a first approximation (see Ref. [64] on thermalization and interaction of DM with OM using the standard cosmological model). The effect of DM on the macroscopic and microscopic properties of the PNS is then investigated using the two-fluid formalism, in which DM and nuclear matter interact only gravitationally. This formalism allows us to introduce a fixed DM mass fraction into stellar matter and investigate its effect.
The OM is composed of nucleons, hyperons, and leptons while the DM, on the other hand, comprises self-interacting dark fermions of the same kind (dark protons and dark neutrons). We use the relativistic mean-field approximation adjusted by the density-dependent meson-nucleon couplings (DDME2) [65] for OM and DM species at $\beta$-equilibrium. The coupling for the hyperons is taken from  [66], which extends the DDME2 to include hyperons and $\Delta$-resonances couplings based on symmetry group arguments. The main focus of this work is to examine how the presence of DM influences temperature distributions, particle distributions, and sound speed within the star throughout its evolution, assuming gravitational interactions. The structure of the resulting star is modeled, and its tidal deformability is calculated to compare the results with observable constraints.
@@ II.1 Nuclear Matter Model
From the matter EoS, i.e. the $P_{OM}$ vs. $\varepsilon_{OM}$ relation, one can obtain the free energy relation, $\mathcal{F}_{B}=\varepsilon_{OM}-Ts$, which allows us to obtain the entropy density $s$ (where $s=s_{B}n_{B}$, with $s_{B}$ the entropy per baryon and $n_{B}$ the total baryon density). The quantities $P_{OM}$ and $\varepsilon_{OM}$ are the respective total pressure and total energy density of the ordinary matter, encompassing all the baryons and leptons present in the system. For stellar matter, it reads
$sT=\varepsilon_{OM}+P_{OM}-\sum_{b}\mu_{b}n_{b}-\sum_{L}\mu_{L}n_{L},$ (7)
@@
$sT=P_{OM}+\varepsilon_{OM}-n_{B}\mu_{B}-\mu_{\nu_{e}}(n_{\nu_{e}}+n_{e}).$ (9)
Equations (8) and (9) are the simplified versions of the $s_{B}$ expression using $\beta$-equilibrium and charge neutrality conditions. We calculate the temperature distributions along with the EoS by fixing $s_{B}$ together with the lepton fraction ($Y_{L,e}$) to control the neutrino concentration in the neutrino-trapped stage and $Y_{\nu_{e}}=0$ in the neutrino-free stage. A detailed calculation of the EoS can be found in [76] where we adopted the EoS for this work. In addition to that, density-dependent meson-nucleon coupling has been used to study PNSs with exotic baryons in [77, 57, 60, 78] where EoSs have been explicitly derived.
PNSs are typically investigated at a fixed $s_{B}$ instead of $T$ because entropy provides a natural framework for describing the thermodynamic state of matter during evolution and ensures computational consistency in studying these objects [58]. During core collapse and bounce, the process is approximately adiabatic outside the shock, leading to a more uniform entropy distribution in different mass shells, which helps characterize the star’s thermodynamic state. Additionally, neutrino emission, which dominates the cooling processes of PNSs, depends strongly on $s_{B}$ and lepton fraction rather than temperature [79, 80]. In contrast, temperature varies significantly with density in newborn NSs, with estimates suggesting it can reach 30–50 MeV in the core while the outer layers remain much cooler. Fixing entropy also facilitates comparisons with supernova simulations, which primarily track entropy evolution rather than temperature [81].
@@ II.2 Dark Matter Model
${\cal L}_{\rm DM}={}$ ${\overline{\psi}_{D}}[(i\gamma_{\mu}\partial^{\mu}-\gamma^{0}g_{v}V_{0})-(m_{D}-g_{\tilde{\sigma}}\tilde{\sigma}_{0})]\psi_{D}$
-- $\frac{1}{2}m_{\rm\tilde{\sigma}}^{2}\tilde{\sigma}_{0}^{2}+\frac{1}{2}m_{v}^{2}V_{0}^{2},$ (10)
@@
This Lagrangian density is similar to the Walecka model for nucleon (see Ref. [82] and references therein), where we have omitted the isospin projection that differentiates between proton and neutron, thus, referring to dark protons and dark neutrons as identical dark fermions. To calculate the EoS for the model, we use the mean-field approximation with density-dependent coupling similar to the approach presented in Sec. II.1 for the OM. Therefore, we replaced $g_{\tilde{\sigma}}$ and $g_{v}$ with $g_{\sigma N}$ and $g_{\omega N}$ and the meson masses $m_{\rm\tilde{\sigma}}$ and $m_{v}$ with $m_{\sigma}$ and $m_{\omega}$ respectively, following Tab. 2, and $m_{D}=939$ MeV. This formalism mirrors the visible sector by adopting the same saturation density, coupling constants, and meson masses as nuclear matter. This assumption reduces the arbitrariness in fixing the couplings. The coupling is defined in the same form as Eq. (5), with $\eta\rightarrow n_{B^{\prime}}/n_{0}$ ($n_{B^{\prime}}$ is the total dark baryon density), using the fit parameters on Tab. 2. The detailed calculation of the EoS of the above Lagrangian density at $T=0$ can be found in [33].
@@
$\varepsilon_{D}={}$ $\sum_{D}\left(\gamma_{D}\int\frac{d^{3}k_{D}}{(2\pi)^{3}}E_{D}\left[f_{D+}+f_{D-}\right]+\frac{g_{\rm\omega N}^{2}}{2m_{\rm\omega}^{2}}n_{D}^{2}\right)$
@@
$n_{D}=\sum_{D}\gamma_{D}\int\frac{d^{3}k_{D}}{(2\pi)^{3}}\left[f_{D\,+}-f_{D\,-}\right],$ (16)
@@
$n_{D}^{s}=\sum_{D}\gamma_{D}\int\frac{d^{3}k_{D}}{(2\pi)^{3}}\frac{m^{\ast}_{D}}{E_{D}}\left[f_{D\,+}+f_{D\,-}\right].$ (17)
@@
$\mu_{p^{\prime},n^{\prime}}^{\ast}=\mu_{p^{\prime},n^{\prime}}-g_{\omega N}V_{0}-\Sigma^{r}_{D},$ (19)
where $\mu_{p^{\prime},n^{\prime}}$ is the real chemical potential of the dark protons ($p^{\prime}$) and dark neutrons ($n^{\prime}$) and $T_{D}$ is temperature. We calculate the EoS at a fixed dark entropy $\tilde{s}$ ($\tilde{s}=s_{D}n_{D}$, where $\tilde{s}$ is DM entropy density, $s_{D}$ is the total entropy per DM particle) using the free energy expression $\mathcal{F}_{D}=\varepsilon_{D}-T_{D}\tilde{s}$ for dark particles. Similarly to the visible sector, the dark sector is treated as a single fermionic fluid governed by the Pauli exclusion principle. Therefore, a balance between dark fermions and dark leptons must be maintained through a dark $\beta$-equilibrium reaction to accurately characterize the physical properties of the dark stellar system. Since we have identical dark fermions, in the neutrino-trapped matter
$\mu_{B^{\prime}}=\mu_{p^{\prime}}+\mu_{e^{\prime}}-\mu_{\nu_{e}^{\prime}},\quad{\rm with}\quad\mu_{n^{\prime}}=\mu_{B^{\prime}}$ (20)
where $\mu_{e^{\prime}}$ is the chemical potential of the dark electrons and $\mu_{\nu_{e}^{\prime}}$ is the chemical potential of the dark electron neutrinos, and in the neutrino transparent matter
$\mu_{B^{\prime}}=\mu_{p^{\prime}}+\mu_{e^{\prime}},$ (21)
with $\mu_{B^{\prime}}$ the total baryon density of the dark sector. The dark muons become relevant when the matter is neutrino-transparent as mentioned above, however, their chemical potential is the same as that of electrons ($\mu_{e^{\prime}}=\mu_{\mu^{\prime}}$). In the mirrored model scenario, DM in isolation can form compact objects (a dark NS, perhaps), with a simpler composition, assuming the presence of only dark nucleons. Since the dark NSs are physically observable objects in the dark world, we impose dark charge neutrality like in the visible sector:
$n_{p^{\prime}}=n_{e^{\prime}},$ (22)
@@
$\tilde{s}T_{D}=P_{D}+\varepsilon_{D}-n_{B^{\prime}}\mu_{B^{\prime}},$ (24)
@@
$\tilde{s}T_{D}=P_{D}+\varepsilon_{D}-n_{B^{\prime}}\mu_{B^{\prime}}-\mu_{\nu_{e}^{\prime}}(n_{\nu_{e}^{\prime}}+n_{e^{\prime}}),$ (25)
in neutrino-trapped matter. Studies related to mirror dark matter models and their applications to compact astrophysical objects can be found in [83, 84, 85, 86].
@@ III Two Fluid Formalism
$\frac{dP_{OM}}{dr}={}$ $-(P_{OM}+\varepsilon_{OM})\frac{4\pi r^{3}(P_{OM}+P_{D})+M(r)}{r(r-2M(r))},$ (26)
$\frac{dP_{D}}{dr}={}$ $-(P_{D}+\varepsilon_{D})\frac{4\pi r^{3}(P_{OM}+P_{D})+M(r)}{r(r-2M(r))},$ (27)
@@
$\frac{dM(r)}{dr}=4\pi(\varepsilon_{OM}+\varepsilon_{D})r^{2}.$ (28)
@@
$F(r)=\frac{r-4\pi r^{3}\left\{[\varepsilon_{OM}(r)+\varepsilon_{D}(r)]-[P_{OM}(r)+P_{D}(r)]\right\}}{r-2M(r)},$ (32)
@@
$Q(r)={}$ $\frac{4\pi r\left\{5[\varepsilon_{OM}(r)+\varepsilon_{D}(r)]+9[P_{OM}(r)+P_{D}(r)]+\frac{\varepsilon_{OM}(r)+P_{OM}(r)}{\partial P_{OM}(r)/\partial\varepsilon_{OM}(r)}+\frac{\varepsilon_{D}(r)+P_{D}(r)}{\partial P_{D}(r)/\partial\varepsilon_{D}(r)}-\frac{6}{4\pi r^{2}}\right\}}{r-2M(r)}$
$-4\left\{\frac{M(r)+4\pi r^{3}[P_{OM}(r)+P_{D}(r)]}{r^{2}\left(1-2M(r)/r\right)}\right\}^{2},$ (33)
@@ IV Results
Table 3: Maximum stellar mass ($M_{\rm max}$), its radius ($R_{\rm max}$), the radii of the canonical and $\rm M=2.1M_{\odot}$ stars ($\rm R_{1.4}$ and $\rm R_{2.1}$), the canonical star tidal deformability ($\Lambda_{1.4}$), and baryonic mass ($M_{\rm B}$) for different configurations, OM compositions, and proportions of DM mass fraction in an NS.
Configuration Composition DM [%] $M_{\rm max}$ [$\rm M_{\odot}$] $R_{\rm max}$ [km] $R_{1.4}$ [km] $R_{2.1}$ [km] $\Lambda_{1.4}$ $M_{\rm B_{2.10}}$ [$\rm M_{\odot}$]
$s_{B}=1$; $Y_{L,e}=0.4$ N 0 2.43 12.49 15.41 14.15 865 2.33
1 2.40 12.40 15.37 14.03 704 2.28
5 2.30 12.12 14.85 13.47 480 2.23
NH 0 2.32 12.53 15.41 14.04 863 2.33
1 2.29 12.46 15.37 13.15 708 2.28
5 2.19 12.12 14.86 13.88 483 2.23
$s_{B}=2$; $Y_{L,e}=0.2$ N 0 2.48 13.01 16.62 15.03 1281 2.34
1 2.45 12.93 16.57 14.89 903 2.29
5 2.35 12.64 15.96 14.50 596 2.25
NH 0 2.30 12.72 16.43 14.42 863 2.35
1 2.28 12.65 16.36 14.24 708 2.29
5 2.18 12.31 15.74 13.08 483 2.25
$s_{B}=2$; $Y_{\nu_{e}}=0.0$ N 0 2.49 13.07 16.97 15.18 1299 2.35
1 2.46 13.02 16.95 15.05 1036 2.27
5 2.34 12.71 16.31 13.82 753 2.25
NH 0 2.28 12.70 16.67 14.36 1152 2.37
1 2.25 12.65 16.63 14.15 988 2.29
5 2.14 12.33 15.96 13.13 737 2.27
$T=0$ N 0 2.48 12.06 13.19 13.20 704 2.44
1 2.45 11.94 13.11 13.08 500 2.40
5 2.33 11.66 12.76 12.61 356 2.36
NH 0 2.26 11.94 13.19 12.98 704 2.35
1 2.22 11.82 13.12 12.80 490 2.27
5 2.11 11.48 12.75 11.83 348 2.24
Figure 1: EoS for dark matter configurations with and without trapped neutrinos (top), and the corresponding temperature distributions (bottom).
@@
Figure 2: The pressure profile of $2.1M_{\odot}$ stellar configuration, considering nucleonic (top) and hyperonic (bottom) admix DM.
Determining the microscopic properties of DANSs, such as particle fractions and temperature distributions, as well as macroscopic thermodynamic quantities like the speed of sound, is inherently complex. This complexity arises from the need to account for gravitational effects, which can only be properly incorporated through the two-fluid TOV equations. In this work, we employ numerical techniques that help us to estimate these quantities without neglecting the gravitational effect. Considering that the stellar configurations are in hydrostatic equilibrium, satisfying the two-fluid TOV equations, the pressure gradient balances the gravitational force at each point inside the star. Thus, the pressure at each point is the sum of the pressures from the two fluids: $P=P_{OM}+P_{D}$. The total mass of the star on the other hand, is determined by integrating over the total energy density, which is the sum of the energy densities of the two fluids: $\varepsilon=\varepsilon_{OM}+\varepsilon_{D}$, as represented in the TOV equations (27) to (28). This approach is particularly useful because the total pressure profile $P(r)$ is derived from the two-fluid TOV equations, which naturally incorporate the effects of gravity on the net pressure and central energy density distribution. The relationship between $P$, $P_{OM}$, and normalized $R$ is shown in Fig. 2. The plot shows that the emergence of new degrees of freedom and increased DM mass fraction increase the pressure in the stellar core.
To estimate the quantities, we calculate the pressure profile of a star with a specific gravitational mass, in this case $2.10M_{\odot}$. The pressure profile is determined by numerically integrating the two-fluid TOV equations from the star’s center ($r=0$) to its surface ($r=R$). Then we extract the pressure, central energy density, and radius corresponding to this star. The integration is performed using the boundary conditions: $P(r=0)=P_{\rm center}$, $M(r=0)=0$, $P(r=R)=0$ and $M(R)=2.10M_{\odot}$ within the two-fluid TOV framework. For each particle fraction or temperature distribution corresponding to $P_{OM}$, we determine its equivalent in $P$ using the cubic spline interpolation method. This serves as a reasonable approximation in the absence of a rigorous first-principles approach that fully accounts for gravity.
@@
$i$ represents the different particles in the stellar matter and $n$ is their number densities. Fig. 3 shows the particle distributions within a star with mass $2.1M_{\odot}$ at different stages of the stellar evolution. Along the panels, we present snapshots of the stellar evolution from core birth to the stages where the star heats up by neutrino diffusion and then to ‘maturity’ as cold-catalyzed NS several years after the supernova explosion. The first panel (top-left) represents the ambient conditions of core birth when the star is neutrino-rich. The second panel (top-right) represents the period of deleptonization after the supernova explosion, at this stage the star is about 0.5 to 1 second old. Here, the stellar matter heats up due to neutrino diffusion, so its temperature is higher than in the first stage. The third stage (bottom-left) occurs when neutrinos fully escape from the stellar core. At this stage, the star reaches its maximum temperature and then begins to cool, roughly 10 to 15 seconds after the core birth. The last panel (bottom-right) is when the star has cooled down through neutrino diffusion and thermal radiation to form a cold-catalyzed NS, over several years. It is important to mention that two fates await the PNS after the supernova explosion: If the pressure generated by the supernova explosion is insufficient to eject the outer mantle of the progenitor massive star completely, the PNS may continue to accrete matter, potentially leading to the formation of a black hole. In this case, it is believed that the neutrino emission can abruptly stop [88] and the evolution of the PNS terminates without moving to the second stage. On the other hand, if the pressure generated by the explosion is sufficient to lift the surrounding stellar envelope and induce deleptonization, the outer mantle collapses, allowing the PNS to continue its evolution. During deleptonization, transitioning from the second to the third panel, the likelihood of strangeness-rich particles emerging in the stellar matter increases. This can abruptly soften the EoS, potentially forming a black hole if the gravitational pressure is sufficiently strong [61]. Therefore, Fig. 3 is constructed assuming that the PNS continues to evolve without collapsing into a black hole.